# Policy Design for Two-sided Platforms with Participation Dynamics

### v1

@@ 7 Real-data Experiment
Simulation.   We simulate the subgroup of viewers and providers, following the procedure presented in (Bose et al., 2023). Specifically, we cluster viewers and providers into $K=L=20$ subgroups respectively, based on the viewer and provider embeddings learned by the neural CF model (He et al., 2017). We use the same initialization and dynamics of the population as described in Section 6. Then, we simulate the utility and population effects as follows.
@@
    Let $u_{k}:=\mathbb{E}[u\,|\,u\in\mathcal{U}_{k}]$ be the mean embeddings of viewer group $k$. We define the group-wise base utility as $b_{l,k}=\mathbb{E}[b(u_{k},c)\,|\,c\in\mathcal{C}_{l}]$ (i.e., mean utility that viewer group $k$ receives providers in group $l$).
@@
    Next, to simulate a population effect, we generate a random permutation of providers within each provider group. Given the current provider population $\lambda_{l}$, we let first $\lambda_{l}$ samples in the permutation as the the set of providers in the subgroup $l$ used at timestep $t$. We denote this subset as $\mathcal{C}_{t,l}$. Then, we define the utility from the provider group $l$ as $q(u_{k},c_{l}):={\arg\max}_{c\in\mathcal{C}_{t,l}}b(u_{k},c)$. Therefore, the population effects are defined as
    $f_{k,l}(\lambda_{l}):={\arg\max}_{c\in\mathcal{C}_{t,l}(\lambda_{l})}b(u_{k},c)-b_{k,l},$
    which increases monotonically as provider population ($\lambda_{l}$) increases.
@@ 8 Related work
This section summarizes the important related work.
@@
Policy optimization under the population departure.  The most relevant existing works to ours are Mladenov et al. (2020) and Huttenlocher et al. (2023), which consider the population dynamics by modeling the departure of viewers and providers. Specifically, Mladenov et al. (2020) assume that a provider will leave the platform if the provider cannot receive adequate exposure (i.e., exposure is below some given threshold). Then, Mladenov et al. (2020) solves the constrained optimization problem as linear integer programming and demonstrates that provider fairness is crucial to maintain a high viewer welfare. To extend, Huttenlocher et al. (2023) additionally consider the departure of viewers who receive less utility than given thresholds. Huttenlocher et al. (2023) also formulate a matching problem to determine which viewers and providers to keep in the platform to achieve high long-term social welfare. However, both works ignore the possible growth of the platform, and how a policy design affects the “growing-the-pie” behavior has remained underexplored. Our work complements these existing work by finding that provider fairness is important to ensure high “population effects” in a generalized formulation.
@@
Policy optimization under strategic content providers.  Another related literature is the policy optimization under strategic content providers (Hron et al., 2022; Jagadeesan et al., 2022; Yao et al., 2023). These works often formulate content providers as “selfish” agents who maximize only their own utility defined by the amount of exposure minus the cost of content generation. As described in Section 2.2, our problem setting can also be seen as a variant of policy optimization under strategic viewers and content providers. However, our formulation is distinctive in modeling the increase and decrease of the total population, while existing works assume that the total number of viewers and providers are fixed. This difference results in novel findings: while Hron et al. (2022) find that more explorative (i.e., stochastic) policy can be against producing high-quality “niche” contents when the population is fixed, provider-fair policy (or stochastic allocation) can be beneficial when taking the population growth of multiple groups into account.
@@
Provider fairness, provider diversity, and viewer welfare. Fairness and diversity among providers have been considered as necessary metrics or constraints when optimizing policies in two-sided platforms (Singh & Joachims, 2018; Wang & Joachims, 2021; Boutilier et al., 2023). While provider-fairness is initially considered important from provider-side perspectives (Singh & Joachims, 2018), recent works considers the impacts of provider fairness on viewer welfare. Specifically, provider fairness turned out important to maintain provider diversity (Yao et al., 2023; Hron et al., 2022), and provider diversity helps maintain viewer welfare in the long-run (Su et al., 2023; Mladenov et al., 2020). Our findings align with these works in pointing out that provider-fairness is important for long-term viewer satisfaction, but from a different viewpoint, suggesting that “population effect” should also be taken into account to invest future growth of populations.
@@ 9 Conclusion
This paper studied the policy design in two-sided platforms where viewer and provider populations matter. Through the control- and game-theoretic analyses, we found that the myopic-greedy policy is guaranteed optimal only when the population effects are linear and homogeneous among provider groups, and otherwise may fall short by ignoring the “population effects”. To take such long-term effects into account, we proposed a simple algorithm that guarantees both viewer satisfaction and provider exposure for future population growth. We believe our work provides a cornerstone to build dynamics-aware allocation policies in two-sided platforms where multiple stakeholders engage.
@@ Acknowledgements
This work was partly funded by NSF CCF 2312774, NSF OAC-2311521, and a gift to the LinkedIn-Cornell Bowers CIS Strategic Partnership. Haruka Kiyohara is supported by the Funai Overseas Scholarship.
We thank Thorsten Jaochims and Richa Rastogi for the early-stage discussion about the dynamics in two-sided platforms.
@@ Appendix B Omitted Proofs
This section provides proofs for the Theorems and Propositions presented in the main text. Note that we define the fixed point and Nash equilibrium as follows.
@@
$\bm{\lambda}_{eq}$ is the fixed point under the policy $\bm{\pi}$ when $\bm{\lambda}_{eq}$ satisfies,
$\exists T_{0}>0,\,\forall t>T_{0},\quad\lVert\bm{\lambda}_{t+1}-\bm{\lambda}_{eq}\rVert\,\leq\,\lVert\bm{\lambda}_{t}-\bm{\lambda}_{eq}\rVert.$
@@ Appendix C Additional experiment settings and observations
Difference of population effects in the synthetic and real-world experiments.   Figure 6 shows the population effect used in the synthetic experiment, defined by Eq. (13). The biggest difference between the synthetic and real-world experiment setting is that we observe saturation of the population effects as an early stage of the provider population growth, i.e., around $\lambda_{l}=100$, which is also a reasonable phenomenon in real-world situations. Therefore, in the synthetic experiment, it is important to distribute the content exposure among multiple subgroups to receive high population effects in many different provider groups. Thus, even the uniform policy outperforms the myopic-policy in this setting. In contrast, when using KuaiRec dataset (Gao et al., 2022), the situation is milder than the synthetic experiment, and therefore the myopic policy works well in the real-world experiment. Together, our synthetic and real-world experiments show that the proposed look-ahead policy performs reasonably well in two different configurations. This is because seeking for both (immediate) viewer utility ($s_{k}$) and provider exposure ($e_{l}$) is important to maximize the look-ahead objective, which depends on reference populations $\bar{\lambda}(s_{k})$ and $\bar{\lambda}(e_{l})$.

### v2

@@ 7 Real-data Experiment
Simulation.   We simulate the subgroup of viewers and providers, following the procedure presented in Bose et al. (2023). Specifically, we cluster viewers and providers into $K=L=20$ subgroups respectively, based on the viewer and provider embeddings learned by the neural CF model (He et al., 2017). We use the same initialization and dynamics of the population as described in Section 6. Then, we simulate the utility and population effects as follows.
@@
    Let $u_{k}:=\mathbb{E}[u\,|\,u\in\mathcal{U}_{k}]$ be the mean embeddings of viewer group $k$. We define the group-wise base utility as $b_{l,k}=\mathbb{E}[b(u_{k},c)\,|\,c\in\mathcal{C}_{l}]$ (i.e., mean utility that viewer group $k$ receives from providers in group $l$).
@@
    Next, to simulate a population effect, we generate a random permutation of providers within each provider group. Given the current provider population $\lambda_{t,l}$, we let first $\lambda_{t,l}$ samples in the permutation as the the set of providers in the subgroup $l$ used at timestep $t$. We denote this subset as $\mathcal{C}_{t,l}$. Then, we define the utility from the provider group $l$ as $q(u_{k},c_{l}):={\arg\max}_{c\in\mathcal{C}_{t,l}}b(u_{k},c)$. Therefore, the population effects are defined as
    $f_{k,l}(\lambda_{l}):={\arg\max}_{c\in\mathcal{C}_{t,l}(\lambda_{t,l})}b(u_{k},c)-b_{k,l},$
    which increases monotonically as provider population ($\lambda_{t,l}$) increases.
@@ 8 Related work
Population shifts.  The most relevant existing works to ours are Mladenov et al. (2020) and Huttenlocher et al. (2023), which consider the population dynamics by modeling the departure of viewers and providers. Specifically, Mladenov et al. (2020) assume that a provider will leave the platform if the provider cannot receive adequate exposure (i.e., exposure is below some given threshold). Then, Mladenov et al. (2020) solve the constrained optimization problem as a linear integer program and demonstrate that provider fairness is crucial to maintain a high viewer welfare. To extend, Huttenlocher et al. (2023) additionally consider the departure of viewers who receive less utility than given thresholds. Huttenlocher et al. (2023) also formulate a matching problem to determine which viewers and providers to keep in the platform to achieve high long-term social welfare. However, both works ignore the possible growth of the platform, and how a policy design affects the “growing-the-pie” behavior has remained underexplored (also, these existing works are not directly applicable to our problem setting). Our work complements these works by finding that provider fairness is important to ensure high “population effects” in a generalized setting.
Note that our work is also closely related to the performative prediction literature (Perdomo et al., 2020; Brown et al., 2022; Narang et al., 2023). This line of work studies the policy optimization under a stationary point of dynamics, when the algorithm affects the states (e.g., population) of the environment, and the optimal algorithm can change depending on the states. While these works (theoretically) discuss the model optimization and its convergence under a convex loss function (Brown et al., 2020), our work empirically demonstrates that the look-ahead policy gradient approach (i.e., non-convex optimization) can work in a complex and interactive dynamics of two-sided platforms.
@@
Strategic content providers.  Another related literature is the policy optimization under strategic content providers (Hron et al., 2022; Jagadeesan et al., 2022; Yao et al., 2023b). These works often formulate content providers as “selfish” agents who maximize only their own utility defined by the amount of exposure minus the cost of content generation. As described in Section 2.2, our problem setting can also be seen as a variant of policy optimization under strategic viewers and content providers. However, our formulation is distinctive in modeling the increase and decrease of the total population, while existing works assume that the total number of viewers and providers are fixed. This difference results in novel findings: while Hron et al. (2022) find that more exploration (randomness) can reduce incentives for producing high-quality “niche” content, we find that a more random provider-fair policy can be beneficial when taking the population growth of multiple groups into account.
@@
Fairness, diversity, and welfare. Fairness and diversity among providers have been considered as necessary metrics or constraints when optimizing policies in two-sided platforms (Singh & Joachims, 2018; Wang & Joachims, 2021; Boutilier et al., 2023). While provider-fairness is initially considered important from provider-side perspectives (Singh & Joachims, 2018), recent works considers the impacts of provider fairness on viewer welfare. Specifically, provider fairness turned out important to maintain provider diversity (Yao et al., 2023b; Hron et al., 2022), and provider diversity helps maintain viewer welfare in the long-run (Su et al., 2023; Mladenov et al., 2020). Our findings align with these works in pointing out that provider-fairness is important for long-term viewer satisfaction, but from a different viewpoint based on the growth of populations.
@@ 9 Conclusion
This paper studies recommender policy design in two-sided platforms where viewer and provider populations matter. Through the control- and game-theoretic analyses, we found that the myopic-greedy policy is guaranteed optimal only when the population effects are linear and homogeneous among provider groups, and otherwise may fall short by ignoring the “population effects”. To take such long-term effects into account, we proposed a simple algorithm and empirically demonstrate that it provides both viewer satisfaction and provider exposure for future population growth. We believe our work provides a cornerstone to build dynamics-aware allocation policies in two-sided platforms where multiple stakeholders engage.
@@ Acknowledgements
This work was partly funded by NSF CCF 2312774, NSF OAC-2311521, and a gift to the LinkedIn-Cornell Bowers CIS Strategic Partnership. The Funai Overseas Scholarship supports Haruka Kiyohara.
We thank Thorsten Jaochims and Richa Rastogi for their thoughtful feedback on the early-stage discussion about the dynamics in two-sided platforms. We also thank anonymous reviewers for the valuable discussion during the review.
Impact Statement
The real-world experiments are run on a public recommendation dataset, and we do not consider there are ethical concerns in our experiments.
For the broader social impact, this paper has shed lights on the population effects in two-sided platforms for the first time. Our findings suggest that ensuring the provider exposure among multiple subgroups, instead of concentrating on a single subgroup, is important to maintain a high population growth and utility. This is also relevant to provider-side fairness (Singh & Joachims, 2018) and diversity (Su et al., 2023) in many decision-making scenarios (Extended related work in Appendix 8 provides a further discussion). Moreover, the population effects potentially matter in broad applications including recommender systems, social networking services, and even school admission or hiring processes. Working on such application-specific configurations would also be an impactful future direction of the research community.
@@ Appendix B Omitted Proofs
This section provides proofs for the Theorems and Propositions presented in the main text. Note that we define the fixed point, stability, and Nash equilibrium as follows.
@@
Let $S(\cdot)$ be the dynamics function that maps the population from the previous timestep to the next timestep, i.e., $\lambda_{t+1}=S(\lambda_{t})$. Then, $\bm{\lambda}_{eq}$ is a fixed point under the policy $\bm{\pi}$ when $\bm{\lambda}_{eq}$ satisfies,
$\bm{\lambda}_{eq}=S(\bm{\lambda}_{eq}),$
where $S$ is a static function under a static policy $\bm{\pi}$.
Definition 2 (Stability).
A fixed point $\bm{\lambda}_{eq}$ is stable under the policy $\bm{\pi}$ when $\bm{\lambda}_{eq}$ satisfies,
$\forall\epsilon>0,\,\exists\delta>0,\quad\lVert\bm{\lambda}_{0}-\bm{\lambda}_{eq}\rVert<\delta\,\implies\,\lVert\bm{\lambda}_{t}-\bm{\lambda}_{eq}\rVert<\epsilon.$
This condition is satisfied when $\mathrm{det}(\nabla_{\lambda}S)<1$ holds.
@@ Appendix C Additional experiment settings and observations
Difference of population effects in the synthetic and real-world experiments.   Figure 6 shows the population effect used in the synthetic experiment, defined by Eq. (13). The biggest difference between the synthetic and real-world experiment setting is that we observe saturation of the population effects as an early stage of the provider population growth, i.e., around $\lambda_{l}=100$, which is also a reasonable phenomenon in real-world situations. Therefore, in the synthetic experiment, it is important to distribute the content exposure among multiple subgroups to receive high population effects in many different provider groups. Thus, even the uniform policy outperforms the myopic-policy in this setting. [^6] Throughout the experiments, we found that there are tradeoffs in concentrating and distributing exposures, and which is better often depends on the problem instance. For example, suppose the special case where viewers do not change their population ($\eta_{k}=0,\,\forall k\in[K]$) and the total viewer population is fixed to 100. Then, consider a scenario with 100 provider groups. In this situation, distributing exposure among different subgroups can result in expected exposure of 1 for each provider group. In such cases, concentrating the exposure to one provider group can be a better strategy than distributing allocation for the total population growth. This is why myopic policy performs well in some scenarios, and our argument is that the proposed method can work adaptively well (i.e., at least better or competitive than both myopic and uniform) regardless if the myopic policy succeeds or falls short. In contrast, when using KuaiRec dataset (Gao et al., 2022), the situation is milder than the synthetic experiment, and therefore the myopic policy works well in the real-world experiment. Together, our synthetic and real-world experiments show that the proposed look-ahead policy performs reasonably well in two different configurations. This is because seeking for both (immediate) viewer utility ($s_{k}$) and provider exposure ($e_{l}$) is important to maximize the look-ahead objective, which depends on reference populations $\bar{\lambda}(s_{k})$ and $\bar{\lambda}(e_{l})$.